A total of $40,000 is to be invested, some in bonds and some in Certificates of Deposits (CDs). If the amount invested in bonds is to exceed that in CDs by $4000, how much will be invested in each type of investment?
Answer by mananth(16946)

Bonds ------------- $x
CD's---------------$y
x+y=40000......................(1)
x-y=4000.......................(2)

Add
2x=44000
/2
x=22000
$22,000 investment in bonds
Balance in CD's